OpenStudy (anonymous):

integral of dx/ x+sqrt(1+2x)?

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ dx }{ x+\sqrt{1+2x} }\]

OpenStudy (anonymous):

trig or algebraic?

OpenStudy (cwrw238):

you might try using the substitution u = 1+ 2x

OpenStudy (anonymous):

means algebraic?

OpenStudy (cwrw238):

yes

ganeshie8 (ganeshie8):

u^2 = 1+2x may also work out nicely

OpenStudy (cwrw238):

yes - thats better - what @ganeshie8 said

OpenStudy (anonymous):

\[u=\sqrt{1+2x}\] \[u ^{2}=1+2x\] \[x=\frac{ u ^{2}-1 }{ 2 }\]

OpenStudy (cwrw238):

thats the way to go

OpenStudy (cwrw238):

now find dx

OpenStudy (cwrw238):

in terms of du

OpenStudy (anonymous):

so,\[dx=\frac{ 1 }{ 2} 2u\] dx=u? then \[\frac{ u }{ \frac{ u ^{2-1} }{ 2 } +u}\]

OpenStudy (cwrw238):

no dx = u du

OpenStudy (anonymous):

ya..thats how it is i forgot du...nut am i doing it right in terms of substituting?

OpenStudy (anonymous):

*nut=but

OpenStudy (cwrw238):

yes we have